40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5943     Entry time: Fri Nov 18 08:29:35 2011
Author: Suresh 
Type: Update 
Category: IOO 
Subject: HEPA air-flow effect on WFS. 

[Koji, Suresh]

    We investigated the effect of airflow from the HEPA filters on the PSL beam fluctuation and the resultant noise injected into the WFS loops.   The hint that the WFS are injecting PSL beam jitter into MC mirror motion lies in the MC2_TRANS_PIT and YAW signal's power spectrum shown here.  First, in the blue trace, which shows the spectrum when the WFS loops are off, we see that the WFS1 and WFS2 error signals have a different shape from that of MC2_TRANS.  Since WFS are affected by the PSL beam jitter while the MC2_TRANS_QPD is not, the WFS spectrum contain excess noise, while the MC2_TRANS signals show only the mirror motion.  Next, upon switching on the WFS1 and WFS2 loops, we notice that the MC2_TRANS  spectra acquire the same shape as the WFS spectra.  This shows that the excess noise from the beam jitter has been injected into the MC2 motion, and shows up in the MC2_TRANS spectra.

   To confirm these conclusions we repeated the above measurement with the HEPA fans at 0% (Blue trace), 20% (Red), 30% (Brown) and  100% (Green).   The plots are shown below.  We can see that there is no difference between 0 and 20% levels but beam jitter is visible at 30% HEPA level.  The WFS loops were ON during this time and we can can see the PSL noise injected in to MC2 motion (Green).

WFS_err_HEPA.png

 

The HEPA filter fans are now at 20%.  How can we be sure that they are really working at 20%, since we cannot see any difference between 0 and 20%?

Now that we have this quiet situation, we also investigated the effect (or lack thereof) of switching on the MC2_TRANS loops.  The figure below shows the spectra with all the loops turned off (Blue), with the WFS1 and WFS2  loops turned on (Green)  and with everything turned on (Red).   With the current output matrix, which is the same simple one as the one in this elog, we see some low frequency suppression.  But it also seems to add some noise into the other WFS loops.  I am not sure of this result, due the long duration of this measurement, the seimic noise level may have changed over the course of this measurement.

WFS_err_mc2t_effect.png

As they are not doing any good just now.  I have turned them off by setting the gain in MC2_TRANS PIT and YAW to zero.

 

ELOG V3.1.3-